Angelo jogged 10 times in 30 minutes around a circular track with a radius of 50 m. Find his average speed and average velocity.



1
Expert's answer
2021-05-07T09:16:48-0400

Average speed,


"v = \dfrac{distance}{time}"


"= \dfrac{10\times2\times 3.14\times 0.5}{30min(60s/min)}"


"= 0.02 m\/s"


Average velocity "= 0 m\/s"